# NewsInterview: a Dataset and a Playground to Evaluate LLMs’ Grounding Gap via Informational Interviews

### v1

@@ 8.4 Annotators
We worked with multiple professional journalists throughout the summer who were either colleagues or students who signed up to work with us. They were all volunteering their time and efforts to help with the research.
@@ 8.5 Computational Resources
All experiments were run either with OpenAI resources (we spent a total of \$$\$300$ running simulations) or open source Llama-3.1-70b models. These models were run on the university cluster, which consisted of either 4xA40s or 2xA100s NVIDIA GPUS.
Appendix A Additional Details on Source Prompts
Persona Description Example Responses
Anxious Unsure if they should be doing the interview, often expresses doubt. "I’m not sure if I should be saying this, I should speak to my manager."
Avoidant Brief, deflects questions, avoids detail, and changes subjects. "Actually, one of the main issues was the supply chain, but we’ve sorted that out."
Adversarial Hostile, challenges the interviewer, provides confrontational replies. "Maybe if you did your job well, you’d understand the data. I’m not here to educate you."
Defensive Protects reputation, feels criticized, gives overly detailed explanations. "One area where costs increased was in material prices, which were out of our control."
Straight- forward Clear, direct, and willing to provide detailed information. "Additionally, we ran out funding midway through the project."
Poor Explainer Struggles to explain clearly, rambles, or provides convoluted answers. "Uh, well, I guess the supply chain was part of it, but, uh, that’s only one part of the story…"
Dominating Controls the conversation, gives lengthy or off-topic answers. "Costs were high, but at my suggestion we brought in the best experts worldwide."
Clueless Confused and uncertain, often unsure of the topic. "Oh, right, the delays… yeah, maybe it was the, uh, supply issues? I’m not too sure…"
Persona Persuasion Description Persuasion Examples
Anxious Responds well to empathetic, reassuring, and patient conversations. Encouraging, non-threatening language builds comfort. "I will be as fair as possible.", "I appreciate your honesty.", "If you’re not comfortable now, I can come back later."
Avoidant Prefers non-obtrusive small talk, short questions, and space to respond. Open-ended, light prompts work well. "And that happened when?", "I imagine there’s more to the story.", "Ah I see."
Adversarial Responds to thorough research, persistence, and fact-based questions. Repeated questioning elicits responses. "Our records indicate…", "Just to be clear, are you saying…?", "Earlier you stated…"
Defensive Engages with non-confrontational and validating conversations. Neutral language reduces defensiveness. "I see why you made that choice.", "We can work together.", "It’s understandable."
Straight-forward Prefers direct and transparent conversations. Efficiency and brevity are key. "Let’s get to the solution.", "What were the key points, in your view?"
Poor Explainer Responds well to structured, patient conversations. Simple clarifying questions and validation help communication. "Explain that part again in smaller steps.", "I understand, keep going.", "Take your time."
Dominating Engages when their expertise is acknowledged. Validation and offering control builds rapport. "I’d love your take.", "You have experience, what do you suggest?", "Your insights are valuable."
Clueless Guided, simple questions with firm direction are effective. Breaking down complex topics increases confidence. "Tell me what you’re thinking.", "It’s okay to be unsure.", "Start with something simple."
Discourse Role Definition
Appendix B Discourse Definitions
[…]
Appendix C Data Preprocessing
C.1 Data Used
@@
C.2 Initial Sizes of the Data
  • There are 1,240,112 rows and 7 columns in utterances-2sp.csv.
  • There are 105,848 rows and 4 columns in episodes.csv.
  • There are 3,199,858 rows and 4 columns in utterances.csv.
@@ C.3.1 NPR-Media
LLM-Generated Counterfactual Questions
C.4 Final Sizes of the Data
@@ C.5 LLM Preprocessing Prompt
C.6 Examples of Interviews
@@
TONY COX, host: "Your Good Thing is About to End."
Ms. MABLE JOHN (Singer): "…Is About to End." Right. Right.
@@
TONY COX, host: In the movie, "Ray," I had looked in the credits to see if there were someone who played you…

### v2

@@ 8.4 Annotators
We worked with multiple professional journalists throughout the summer who were either colleagues or students who signed up to work with us. They volunteered their time and efforts to help with the research.
@@ 8.5 Computational Resources
All experiments were run either with OpenAI resources (we spent a total of $\mathdollar 300$ running simulations) or open source Llama-3.1-70b models. These models were run on the university cluster, which consisted of either 4xA40s or 2xA100s Nvidia GPUS.
Acknowledgments
The authors would like to thank Bloomberg for generously supporting this work by funding the first author with a life-changing 4-year PhD fellowship. Funds from the Defense Advanced Research Projects Agency (DARPA) under Agreement No. HR00112490374 also contributed. Any opinions, findings, conclusions, or recommendations expressed here are those of the authors and do not necessarily reflect the view of our sponsors.
Appendix
Appendix A Additional Dataset Details
In this section, we share additional details of our dataset. We provide examples illustrating the differences between human interviewer behavior and LLM-generated counterfactual questions.
The dataset consists of 45,848 transcribed dialogues with speaker annotations that distinguish between hosts and guests. When examining turn alternation patterns by filtering consecutive turns from the same speaker, we observe an average 7.6 distinct speaking turns per interview.
A linguistic analysis uncovers differences in verbosity between speakers. Sources produce an average of 27.7 words per utterance, while interviewers maintain concise interactions at 16.2 words per utterance. This 71% difference in utterance length reflects the distinct communicative roles: interviewers pose targeted questions while sources provide detailed explanations.
The interview corpus spans a diverse range of subject areas. As can be seen in Table 5, running LDA with $k=7$ over our corpus yields topics from literature and storytelling to political discussions, academic/scientific coverage, and international affairs. It also includes conversations on defense and military topics, economic and business news, and legal/crime issues.
The majority of questions posed by interviewers are of the “What/Which” variety (36.2 %), followed closely by “Other” types (35.6 %) and “How/Why” questions (19.6 %), with “When/Where” and “Who” questions comprising only 5% and 3.4%, respectively (Table 3). In terms of linguistic complexity, interviewer utterances tend to be shorter than guest (source) utterances, averaging 16.18 words per utterance and 270.86 words per conversation for interviewers versus 27.69 words per utterance and 551.92 words per conversation for sources (Table 4). Interviewers also use slightly shorter words on average (4.21 characters) compared to sources (4.67 characters), and exhibit marginally higher Flesch Reading Ease (70.9) Kincaid et al. (1975) than sources (68.4), although both groups share an identical Grade Level score (6.9) (Table 4).
Question Type Frequency Percentage
What/Which 19,553 36.2%
Other 19,236 35.6%
How/Why 10,625 19.6%
When/Where 2,719 5%
Who 1,854 3.4%
Table 3: Distribution of Question Types by Interviewers. Question types posed by human interviewers are measured by searching for keywords in each interviewer utterance.
Metric Interviewer Source
Avg. Words / Utt. 16.18 27.69
Avg. Words / Conv. 270.86 551.92
Avg. Word Len 4.21 4.67
Read Ease 70.9 68.4
Grade Level 6.9 6.9
Table 4: Linguistic Metrics by Speaker Type. Reading Ease and Grade Level measured by Kincaid et al. (1975)
Table 6 demonstrates how LLMs tend to generate follow-up questions when humans would transition to new topics, while Table 7 shows LLMs failing to provide acknowledgment statements that humans naturally use to build rapport. Table 8 presents a representative sample of LLM-generated interview questions, showcasing their typical structure and content when prompted to continue human interview conversations.
Literature Politics Academia International Affairs Defense Economy Legal
book vote melissa inskeep military company court
write republican block china kelly market school
author senator norris iran security percent police
life party michele north iraq siegel jacki
play campaign university countries afghanistan bank lyden
story obama audie madeleine force economy justice
game election flatow brand russia business judge
love political cornish nuclear israel price charge
women democrats water international intelligence cost information
read republicans professor korea general program department
Table 5: Top 10 words for each of the $k=7$ LDA topics over the interview corpus. Column headers are 1–2 word summary keywords, and each column is shaded in a different light color.
Prior Context:
“…MELISSA BLOCK, HOST: Charlie, I gather there’s another race in North Carolina that you’ve been paying some attention to.
CHARLIE MAHTESIAN: Of course. I can’t help myself. Remember Clay Aiken, the ‘American Idol’ runner‐up from a few years ago back?
MELISSA BLOCK, HOST: I do. I do.
CHARLIE MAHTESIAN: How can we forget him? Well, he’s running for Congress as a Democrat this year. And it turns out that in addition to being a talented singer, he’s a pretty decent candidate. But despite the great name recognition that he got from his showbiz career, he’s not a lock to win. And in any case, if he does win the Democratic nomination tomorrow, it’s still something of an uphill climb for him because he’s going to be running in a pretty comfortable and safe Republican district.”
Actual Question
Outline-Level Question
MELISSA BLOCK, HOST: Let’s move over to Ohio.
LLM Question
Follow-Up Question
What makes Clay Aiken a decent candidate, and what are his chances of winning in a safe Republican district?
Table 6: Example of an entire LLM counterfactual generation sequence. The prior k=4 turns of human QA-pairs is fed in as context. Next, the human interviewer’s turn is shown, compared with the LLM generation. The LLM used for generation is the baseline prompting LLM. Labels for each question are shown as well.
Prior Context:
“ARUN RATH, HOST: As a father of two young kids, I am very excited about a new segment we’re starting right now about parenting…
HANK AZARIA: I am not a children kind of person. I don’t really like kids. I don’t gravitate towards them. They make me nervous when I’m around them. I didn’t particularly like myself as a child.
ARUN RATH, HOST: Well, Hank did become a father, and his son is now 4. I know from personal experience that’s when kids start to ask deep, complicated questions. I wanted to know how Hank handled one of the big ones: What happens when you die?
HANK AZARIA: Sadly, my dad passed away this past year, so death came up with my son. And it was amazingly hard. He got really upset about it, as you might imagine. I mean, he was like, what do you mean we’re not going to see Grandpa anymore? …
ARUN RATH, HOST: Wow.
HANK AZARIA: And then God gave us a son. How about that for just a shot-in-the-dark theory of how it happens? But, man — and he really is with it. He’s like, hey, can I have a baby someday? And we’re like, sure you can. Of course, you can. He goes, can I borrow your magic wand? And I’m like, you’ll have your own. You’ll have your own. Anyway, just wanted to share that one.”
Actual Question
Acknowledgement Statement
ARUN RATH, HOST: That was great.
LLM Question
Follow-Up Question
How do you think your son’s understanding of death and the concept of making babies has evolved over time, and have you had to revisit these conversations as he’s gotten older?
Table 7: A second example of an LLM counterfactual generation sequence. The human interviewer is giving a significant amount of encouragement and acknowledgment throughout the interview, compared with the LLM. Even when the LLM sees this speech pattern, it does not imitate it.
LLM-Generated Counterfactual Questions
Appendix B Data Preprocessing
B.1 Data Used
@@
B.2 Initial Sizes of the Data
  • There are 1,240,112 rows and seven columns in utterances-2sp.csv.
  • There are 105,848 rows and four columns in episodes.csv.
  • There are 3,199,858 rows and four columns in utterances.csv.
@@ B.3.1 NPR-Media
B.4 Final Sizes of the Data
@@ B.5 LLM Preprocessing Prompt
(a) Clueless
(b) Dominating
(c) Avoidant
Figure 7: Beta Distributions for Various Interview Personas
B.6 Examples of Interviews
@@
TONY COX, host: “Your Good Thing is About to End.”
Ms. MABLE JOHN (Singer): “…Is About to End.” Right. Right.
@@
TONY COX, host: In the movie, “Ray,” I had looked in the credits to see if there were someone who played you…